### v2

# Dislocation correlations in GaN epitaxial films revealed by EBSD and XRD

Abstract
Correlations between dislocations in crystals reduce the elastic energy via screening of the strain by the surrounding dislocations. We study the correlations of threading dislocations in GaN epitaxial films with dislocation densities of $5\times 10^{8}$ cm-2 and $1.8\times 10^{10}$ cm-2 by X-ray diffraction (XRD) in reciprocal space and by high-resolution electron backscatter diffraction (HR-EBSD) in real space, where the strain is derived from a cross-correlation analysis of the Kikuchi patterns. The measured XRD curves and HR-EBSD strain and rotation maps are compared with Monte Carlo simulations within one and the same model for the dislocation distributions. The screening of the dislocation strains is modeled by creating pairs of dislocations with opposite Burgers vectors, with the mean distance between dislocations in a pair equal to the screening distance. The pairs overlap and cannot be distinguished as separate dipoles. The HR-EBSD-measured autocorrelation functions of the strain and rotation components follow the expected logarithmic law for distances smaller than the screening distances and become zero for larger distances, which is confirmed by the Monte Carlo simulations. The kink in the plot of the autocorrelation function allows a robust and accurate determination of the screening distance without making any simulation or fit. Screening distances of 2 µm and 0.3 µm are obtained for the samples with low and high dislocation densities, respectively. The dislocation strain is thus screened by only 4 neighboring dislocations. In addition, an anisotropic resolution of the HR-EBSD measurements is observed and quantified.
Keywords: X-ray diffraction, electron backscatter diffraction, dislocations, strain, GaN
1 Introduction
It is well established that the elastic energy of a single straight dislocation per unit length of the dislocation line $E\propto(\mu b^{2}/4\pi)\ln(R/r_{c})$ diverges as the lateral crystal size $R$ is increased to infinity. Here $\mu$ is the shear modulus, $b$ is the length of the Burgers vector, and $r_{c}$ is the dislocation core radius ($r_{c}\simeq 2.6b$) [1]. Hence, the elastic energy density of a crystal containing a finite density of random uncorrelated dislocations tends to infinity as the crystal size increases. The energy remains finite when dislocations are correlated, such that the long-range strain field of a dislocation is screened by surrounding dislocations. Then, $R$ is the screening distance, rather than the crystal size.
Figure 1: (a,b) Panchromatic cathodoluminescence intensity maps of GaN samples 0 and 1 and (c) scanning electron micrograph of sample 2.
Both dislocation density and the screening distance affect the intensity distributions in X-ray diffraction (XRD). The XRD line profile analysis [2, 3] is an established method in powder diffraction for determining the dislocation density as well as the screening distance, and hence the elastic energy stored in a crystal, and changes to these quantities during plastic deformation. The calculation of the diffracted intensity requires a model for the dislocation arrangement. Wilkens proposed [4, 5] a “restrictedly random dislocation distribution”, where the crystal is subdivided into cells, with each cell containing the same number of dislocations. The numbers of dislocations with opposite Burgers vectors in a cell are equal so that the total Burgers vector of a cell is zero, resulting in the total strain from the dislocations in a cell decaying with the distance faster than the strain of an individual dislocation (in Wilkens’ model, the strain is assumed to be restricted within the cell). The cell size $R$ then plays the role of the screening distance for dislocation strains. The strain probability density distribution calculated for screw dislocations in the Wilkens model [5] is used in computer programs for XRD line profile analysis [6, 7].
The XRD intensity distribution can be calculated for any given dislocation distribution model by means of Monte Carlo integration [8, 9]. A Monte Carlo simulation of the XRD profiles that implements the Wilkens model literally leads to artifacts when a cell contains only one or a few dislocation pairs [10]. An alternative model for dislocation strain screening is to consider random pairs of dislocations with opposite Burgers vectors, with $R$ being the mean distance between dislocations in a pair [11]. When $R$ exceeds the mean distance between all dislocations $\varrho^{-1/2}$, the pairs overlap and cannot be distinguished as individual dipoles. Below, we only deal with such overlapping pairs and do not consider the case of dislocation dipoles with the dipole widths small compared to the separation between dipoles.
In laboratory XRD studies of dislocations, the diffracted intensity is collected from the entire sample and integrated over the diffraction vector orientation in powder diffraction or the diffracted wave direction in a double-crystal diffraction setup for single crystals. The XRD line profile analysis is based on fitting of the experimental curves involving two parameters of the dislocation distribution, the dislocation density $\varrho$ and the screening distance $R$. In many XRD studies, the dislocation density $\varrho$ is of primary interest, while the screening distance $R$ is an auxiliary fit parameter required for accurately determining the dislocation density. However, determination of the elastic energy stored in a crystal requires an accurate determination of $R$ [12]. The screening distance $R$ is accessed as a transition from the Gaussian central part of the diffraction line to its tails, which show an intensity decay that is proportional to $q^{-3}$. These tails originate from diffraction in highly disturbed regions close to dislocation lines and their intensity is independent of dislocation correlations.
The screening distance $R$ can be accessed directly if spatial strain maps are measured. $R$ can then be obtained from the strain-strain correlation function [13]. In that work, all components of the strain and rotation tensors have been measured on the example of threading dislocations in (In,Ga)N films by the scanning XRD microscopy technique [14], using synchrotron X-rays. Mapping in three non-coplanar reflections was performed to obtain all nine components of the strain and rotation tensors. Since the dislocation strain depends on the distance from the dislocation line $r$ by the universal law $\varepsilon\propto r^{-1}$, the correlation function decreases proportionally to the logarithm of the distance for distances smaller than $R$, and becomes zero at larger distances. Plotting the correlation function against the logarithm of the distance yields a hook-shaped curve. The kink in this curve provides the screening distance without the need for any fitting or assumptions about the dislocation correlations.
A variety of lattice-sensitive microscopy techniques for mapping strain and rotation in crystals is nowadays available [15]. A number of synchrotron-based XRD techniques, such as X-ray Laue microdiffraction [16, 17], dark-field X-ray microscopy [18], scanning XRD microscopy [14], Bragg coherent diffraction imaging [19], or Bragg X-ray ptychography [20], provide access to the full strain tensor. The correlations of strain fields have been studied to characterize dislocation distributions by scanning XRD microscopy [13] and dark-field X-ray microscopy [21].
While these XRD techniques are restricted to synchrotron radiation sources, high-resolution electron back scattering diffraction (HR-EBSD) is a widely available laboratory technique that provides spatial maps of all components of the strain and rotation tensors near the crystal surface. The strains and rotations are obtained from shifts of the Kikuchi bands [22, 23, 24]. The strain sensitivity of the method was enhanced to $10^{-4}$ with the invention of the a cross-correlation-based analysis of EBSD patterns [25, 26]. Strain and rotation maps, as well as the strain probability density distributions, have been reported in several studies [27, 28, 29]. The maps obtained in the present study are similar to those reported in the previous studies. However, the primary aim of our work is to obtain and analyze the spatial correlations of the strains, a subject not addressed in these previous studies. We demonstrate that the strain-strain correlations calculated from the maps clearly reveal the dislocation strain screening by surrounding dislocations and provide the screening distances $R$.
The analysis of dislocations in GaN and other group III nitrides is of primary interest for optoelectronic applications of these materials and, at the same time, can serve as a touchstone for various methods of evaluating dislocation arrays. GaN epitaxial films contain threading dislocations oriented parallel to each other and perpendicular to the film surface. These well-defined dislocation arrays enable detailed modeling of both XRD curves and HR-EBSD maps and their quantitative comparison. A respective comparison of different methods for dislocation density determination in metals [30] is much more complicated since it requires an account of several slip systems, grain boundaries, etc.
We study GaN(0001) epitaxial films with dislocation densities differing by orders of magnitude. We compare a film grown by metal-organic chemical vapor deposition with a threading dislocation density of $5\times 10^{8}$ cm-2 and a film grown by molecular beam epitaxy with a dislocation density of $1.8\times 10^{10}$ cm-2. A free standing GaN film with a dislocation density as low as $6\times 10^{5}$ cm-2 and a perfect Si crystal are used as references.
We carry out both laboratory XRD and HR-EBSD measurements on the same samples and also perform Monte Carlo simulations of the XRD line profiles and HR-EBSD strain maps, using one and the same model of dislocation distributions for both experimental techniques and the same computing program. Comparing the measured and Monte-Carlo-modeled strain maps in HR-EBSD provides us with two further findings in addition to the determination of the screening distance for the dislocation strain. First, we find that the resolution of HR-EBSD is highly anisotropic, with significantly worse resolution in the direction of the inclination of the incident electron beam. We obtain a quantitative estimate of the resolution by comparing the strain-strain correlations in two orthogonal directions. Secondly, we find that the experimental HR-EBSD maps show, as uncorrelated random noise, shear strain components that should be zero at a strain-free surface. The widths of their distributions can be used as an estimate of the accuracy of the HR-EBSD measurements.
2 Experiment and Monte Carlo simulations
2.1 Samples
We study three GaN(0001) samples with different densities of threading dislocations (see Fig. 1). In all samples, the dislocations are straight lines running from the substrate to the film surface, along the surface normal. Hence, threading dislocations provide well defined arrays in single crystal samples suitable for both XRD and HR-EBSD studies. The investigated samples were characterized in our previous works [31, 32, 33].
As a reference, we use a 350 µm-thick free-standing GaN (0001) film grown by hydride vapor phase epitaxy. We refer to it as sample 0. The dislocation density in this film is estimated from cathodoluminescence images to be as low as $6\times 10^{5}$ cm-2. Since this sample produced an unexpectedly broad strain distribution in the HR-EBSD measurements described in Sec. 3.2 below, it was rinsed with isopropanol and ethanol and then cleaned with oxygen plasma to remove any contaminants and produce a uniformly oxidized surface layer. Etching of the surface with heated KOH removed this oxidized layer and left a smooth surface with a root-mean-square roughness of 0.31 nm. However, this cleaning step had very little effect on the HR-EBSD maps.
Sample 1 is a 5.6 µm-thick GaN(0001) film on an Al2O3(0001) substrate. A 1.3-µm-thick GaN layer is fabricated by plasma-assisted molecular beam epitaxy on top of a 4.3-µm thick GaN(0001) template, which in turn is grown by metal-organic chemical vapor deposition on an Al2O3(0001) substrate. As this sample has been grown for our former study [32, 33], an additional 3-nm-thick (In,Ga)N single quantum well is buried at about 650 nm below the surface, which however is not relevant for the present work. Threading dislocations, which are in the focus of the present study, are inherited from the template. Figure 1(b) presents a cathodoluminescence (CL) intensity image of this sample. Dislocations are seen as dark spots since they act as centers of nonradiative recombination of excitons [34, 35, 36, 37, 38]. The density of the spots is $5\times 10^{8}$ cm-2. We note that the counting of the dark spots on CL images may underestimate the density of threading dislocations [39].
Sample 2 has been grown for another study of ours [31]. A 2.5 µm-thick GaN(0001) film is grown by plasma-assisted molecular beam epitaxy on a 6H-SiC(0001) substrate. A cathodoluminescence image of this sample shows overlapping dark spots because of the large dislocation density, which does not allow an accurate determination of the dislocation density. Figure 1(c) presents a scanning electron micrograph of this sample. It shows dark spots with the density of $1.8\times 10^{10}$ cm-2. These spots have been identified as pits at dislocation outcrops by comparing a similar electron micrograph of sample 1 with the CL map taken from the same surface area [33]. For the sample with a low dislocation density, a one-to-one correspondence of the pits with the dark spots in the CL image has been observed.
Additionally, a reference HR-EBSD measurement is obtained on a Si(001) wafer.
2.2 X-ray diffraction
[…]
2.3 Electron backscatter diffraction
Scanning electron microscopy (SEM) and CL imaging, as well as the HR-EBSD measurements were carried out in a Zeiss Ultra 55 scanning electron microscope. For the HR-EBSD measurements, the microscope was operated at an acceleration voltage of 15 kV with a beam current of 6 nA. To record the Kikuchi patterns, the stage with the mounted sample was tilted by 70 with respect to the incident beam direction towards an EDAX Hikari Super EBSD detector.
The patterns were recorded with the highest available resolution of $470\times 470$ pixels and background-corrected using a reference pattern. HR-EBSD maps were recorded over either $2\times 2$ µm2 or $10\times 10$ µm2 areas with step sizes of 20 nm and 50 nm, respectively, at an exposure time of 100 ms. Measurements with these settings took approximately 20 min for the smaller maps and 80 min for the bigger maps to record. Sample drift was estimated by comparing SEM images recorded before and after the measurement and were smaller than 18%. Its effect is quantified and discussed at the end of Sec. 3.3, as the discussion is based on the results of Monte Carlo simulations. We find that the effect of drift can be neglected.
Cross-correlation analysis of the Kikuchi patterns [^1] Early studies distinguish Kikuchi patterns formed by electron diffraction in transmission electron microscopy, and pseudo-Kikuchi patterns due to diffraction of backscattered electrons in SEM [43]. Nowadays, the term Kikuchi pattern is used for both of them. [44] was carried out with the software CrossCourt 4.5.3.6 from BLG Vantage. Examples of Kikuchi patterns from all four samples are presented in Fig. SM1 in the Supplementary Material (SM). The software defaults were chosen for 20 regions of interest (squared subsets of the Kikuchi pattern). A reference pattern was chosen as the pixel with the lowest kernel average misorientation in the initial indexation of the EBSD map. The Kikuchi pattern of each region was compared to the same area of a reference pattern to identify tiny shifts and rotations of the pattern. The cross-correlation method enables these changes to be determined with a sub-pixel resolution. From changes in different parts of the Kikuchi pattern, the strain and rotation tensor components are obtained using the elastic moduli of GaN [45]. Applying this analysis to every point of an EBSD map enables the changes in strain and crystal rotation to be mapped.
2.4 Monte Carlo simulations
Monte Carlo simulations of both XRD curves and HR-EBSD maps are performed in the framework of one and the same model of dislocation arrangements and dislocation strains using the computing programs that were developed earlier for XRD [31, 46] and adapted here for HR-EBSD. We include in the simulations all types of perfect threading dislocations that are present in wurtzite GaN epitaxial layers [47, 48, 49, 50]. The displacement fields of a-type edge dislocations with Burgers vectors $1/3\left\langle 11\bar{2}0\right\rangle$, c-type screw dislocations with Burgers vectors $\left\langle 0001\right\rangle$, and a+c-type mixed dislocations with Burgers vectors $1/3\left\langle 11\bar{2}3\right\rangle$ are calculated. Crystallographically equivalent Burgers vectors are taken with equal probabilities.
The screening of the dislocation strains by surrounding dislocations is provided by generating dislocations in pairs with opposite Burgers vectors. The positions of the pairs are random and uncorrelated. The distances between dislocations in the pairs are taken from a lognormal distribution with a mean value of $R$ (hereafter referred to as the screening distance) and a standard deviation of $R/2$. The simulations are insensitive to the standard deviation: a unimodal distribution (all dislocation pairs have the same distance between dislocations) gives the same result. The XRD intensity $I(q)$ is calculated as a probability density of the distortion component $q=\hat{\mathbf{K}}{}^{\mathrm{out}}\cdot\nabla(\mathbf{Q}\cdot\mathbf{U})$, where $\mathbf{U}(\mathbf{r})$ is the total displacement due to all dislocations, $\mathbf{Q}$ is the diffraction vector, and $\hat{\mathbf{K}}{}^{\mathrm{out}}$ is the unit vector in the direction of the diffracted beam [31, 46]. It is worth noting that the direction $\hat{\mathbf{K}}{}^{\mathrm{out}}$ emerges in the diffraction from a single crystal as a consequence of the intensity integration performed by a widely open detector over the plane in reciprocal space perpendicular to $\mathbf{K}^{\mathrm{out}}$. In powder diffraction, the respective intensity integration from randomly oriented grains is performed over the plane perpendicular to the diffraction vector. In this case, $\hat{\mathbf{K}}{}^{\mathrm{out}}$ is replaced with the unit vector in the direction of $\mathbf{Q}$ [42].
The HR-EBSD maps are simulated by calculating all components of the strain tensor $\varepsilon_{ij}=(\partial U_{i}/\partial x_{j}+\partial U_{j}/\partial x_{i})/2$ and the rotation tensor $\omega_{ij}=(\partial U_{i}/\partial x_{j}-\partial U_{j}/\partial x_{i})/2$, using the same displacement vector $\mathbf{U}(\mathbf{r})$. For the XRD simulations, it is sufficient to use the displacement field of a straight dislocation in an infinite crystal, as the X-ray intensity is collected over the entire film thickness. The displacement for a dislocation along the $\left\langle 0001\right\rangle$ direction in a hexagonal crystal is employed. It coincides with the isotropic solution with the Poisson ratio taken to be $\nu_{h}=c_{12}/(c_{11}+c_{12})$ ([51], Sec. 2.5). Here $c_{ij}$ are the elastic moduli. Using their values for GaN [45], we obtain $\nu_{h}=0.27$. For the HR-EBSD simulations, the elastic strain relaxation at the free surface is taken into account. We have compared the isotropic solution ([52], Sec. 6.3) with a more comprehensive anisotropic solution for a dislocation along the sixfold axis ([51], Sec. 2.5). The difference is negligible compared to the statistical errors in the Monte Carlo simulations presented below.
3 Results
3.1 X-ray diffraction
Figure 2: (a,b) X-ray diffraction curves of samples 1 and 2 (thick gray lines) and their Monte Carlo simulations (thin black lines). (c,d) Dislocation density and the screening distance $R$ obtained in the fits of the experimental (full symbols) and the simulated (open symbols) curves. The fits for different reflections are made independently and presented as a function of the angle $\Psi$ between diffraction vector and surface. The parameters for sample 1 are presented by black symbols, and for sample 2 by blue symbols. Dashed lines in (d) mark the values of screening distances obtained in the analysis of EBSD maps.
Figures 2(a) and 2(b) show by thick gray lines measured XRD curves for several reflections from samples 1 and 2, respectively. A symmetric Bragg reflection $0002$ and two asymmetric ones, [^2] We refer to these reflections as asymmetric following the common tradition. In fact, the diffraction vectors of these reflections are inclined with respect to the surface normal, and these reflections are asymmetric in coplanar diffraction. In skew geometry, however, these reflections are symmetric (the incident and the diffracted waves make equal angles with the crystal surface) and non-coplanar (the scattering plane does not contain the surface normal) [42]. $411\bar{2}$ 4 and $022\bar{2}$ 02, are presented. Some qualitative conclusions can be drawn from a direct comparison of the curves. The diffraction vector of the symmetric reflection is directed along the lines of threading dislocations. Consequently, this reflection is only sensitive to screw dislocations and screw components of mixed dislocations, which produce displacements in the direction of the dislocation lines. The widths of the symmetric and asymmetric reflections from sample 1 are closely comparable, indicating similar densities of screw and edge dislocations. In contrast, the symmetric reflection from sample 2 is notably narrower than the asymmetric reflections from this sample, which is due to a low density of screw dislocations in comparison with the edge ones. Comparing diffraction curves from the two samples, one can see that the density of screw dislocations in sample 2 is smaller than in sample 1, while the density of edge dislocations is notably larger.
The thin black lines in Figs. 2(a,b) represent Monte Carlo simulations of the respective diffraction curves. First, we simulate the symmetric Bragg reflections of sample 1 with screw dislocations. Agreement with the experimental diffraction profiles is obtained with a density of $8\times 10^{8}$ cm-2 and the screening distance of $R=1$ µm. Next, we turn to asymmetric reflections and include the edge components of the Burgers vectors in the calculations. Since screw dislocations are very uncommon in GaN(0001) films grown by metal-organic chemical vapor deposition [47, 48, 49, 50], we attribute the density of screw dislocations obtained in the symmetric reflections to the screw components of mixed a+c-type dislocations. The calculation of the diffraction profiles of asymmetric reflections with mixed dislocations agrees with the experimental diffraction curves, meaning that we do not need to add extra edge dislocations. We therefore conclude that sample 1 contains only mixed dislocations.
For sample 2 grown by molecular beam epitaxy, we also first simulate the symmetric Bragg reflections with screw dislocations. This revealed a screw dislocation density of $5\times 10^{8}$ cm-2 and a screening distance of $R=0.7$ µm. These dislocations provide only a small contribution to the widths of the asymmetric reflections. To simulate these reflections, we add edge dislocations with a density of $5\times 10^{10}$ cm-2, two orders of magnitude higher than the density of screw dislocations, with the screening distance of $R=0.3$ µm.
We fit both the measured and the Monte-Carlo-simulated XRD curves in the same way, following Ref. [42]. The symmetric reflections are fitted first to get the densities of screw dislocations, and their contributions are taken into account in the fits of the asymmetric reflections, which give the densities of the edge dislocations. Each reflection is fitted independently from the others. The two essential fitting parameters are the dislocation density $\varrho$ and the screening distance $R$. Figures 2(c,d) present by full symbols the values of $\varrho$ and $R$ obtained in the fits of the measured XRD curves, while open symbols are the results of the corresponding fits of the Monte-Carlo-simulated curves. The data are plotted as a function of the angle $\Psi$ between the diffraction vector and the surface. Thus, $\Psi=0$ corresponds to an in-plane reflection and $\Psi=90^{\circ}$ to a symmetric Bragg reflection. Plotting the fit parameters as a function of $\Psi$ is just a convenient way of presenting the data: ideally, no variation of the parameters is expected.
The error bars representing the possible errors in determining the fit parameters from a fitted curve of a given reflection are smaller than the symbol size in Figs. 2(c,d) and can therefore be neglected. The scattering of the data points obtained from fitting the experimental curves in different reflections can have two sources, the accuracy of the fit function proposed in Ref. [42] and the presence of additional sources of strain, particularly a contribution from dislocations at the interface between the substrate and the relaxed epitaxial films. Our Monte Carlo simulations enable us to estimate the effect of the former by comparing the input parameters of the simulation with the output parameters of the fit.
For sample 1, the input parameters for the Monte Carlo simulation were the dislocation density $\varrho=8\times 10^{8}$ cm-2 of mixed dislocations and the average width of the dislocation pairs $R=1$ µm. Fitting the Monte Carlo simulated curves gives a dislocation density $\varrho=(8.3\pm 0.4)\times 10^{8}$ cm-2 and an average width of the dislocation pairs $R=1.4\pm 0.45$ µm. The errors represent the scattering between data points of different reflections in Figs. 2(c,d). For sample 2, the input parameters for the Monte Carlo simulations were the density $\varrho=5\times 10^{10}$ cm-2 of edge dislocations and $R=0.3$ µm. The fit of the simulated curves gives $\varrho=(5\pm 0.06)\times 10^{10}$ cm-2 and $R=(0.5\pm 0.05)$ µm. These results show that the fits following Ref. [42] quite accurately reproduce the dislocation density, but overestimate the screening distance by about 50%. This is not surprising since the fit formula for XRD curves contains prefactors for the dislocation density and for the screening distance. The prefactor for the dislocation density, also referred to as the “contrast factor” in powder diffraction, depends on the relative orientations of all the vectors involved in the problem (the diffraction vector, the Burgers vector, and the dislocation line direction). It can be accurately calculated in powder diffraction [53] and in diffraction from single crystals [42]. A respective prefactor for the screening distance has only been calculated analytically in the case of screw dislocations in the Wilkens’ model [5]. It can only be estimated for other cases, which limits the accuracy of determining the screening distance from the XRD profiles [12].
Fitting the experimental curves in different reflections yields for sample 1 a density $\varrho=(1\pm 0.3)\times 10^{8}$ cm-2 of mixed dislocations and $R=1\pm 0.7$ µm, and for sample 2 a density $\varrho=(5.8\pm 0.7)\times 10^{10}$ cm-2 of edge dislocations and $R=0.46\pm 0.14$ µm. These dislocation densities are larger than the densities of the spots in the cathodoluminescence or scanning electron microscopy images in Fig. 1 by factors of 2 for sample 1 and 3 for sample 2. We note that similar overestimates of the dislocation density in comparison with transmission electron microscopy results were obtained in previous studies [42, 54, 55].
The HR-EBSD study presented in the next section complements the reciprocal-space analysis by XRD with a real-space observation of the dislocation correlations. In this way, the screening distance $R$ can be determined more accurately and independent of the dislocation density determination. The values of $R$ obtained from the HR-EBSD maps are marked by dashed lines in Fig. 2(d) for comparison.
3.2 Electron backscatter diffraction: experiment
Figure 3: HR-EBSD maps of the strain component $\varepsilon_{22}$ (a–c) in GaN samples 0, 1, and 2 and (d) in a Si(001) wafer, and (e,f) Monte Carlo simulation of the $\varepsilon_{22}$ maps for samples 1 and 2. Note that the strain obtained for Si(001) is an order of magnitude lower than in the GaN samples. The $y$ axis is chosen in the direction of the tilt of the electron beam.
Figures 3(a–c) present maps of the strain component $\varepsilon_{22}$ of the GaN samples 0, 1, and 2 obtained from the analysis of the HR-EBSD measurements. Complete sets of maps of the six components of the strain tensor $\varepsilon_{ij}$ and the three components of the rotation tensor $\omega_{ij}$ are presented in the SM, Figs. SM2, SM3, and SM4. As the HR-EBSD maps are relative to an arbitrarily chosen reference value, we set the mean value of each map to zero. We follow the standard HR-EBSD axes notation: the sample surface is the $xy$ plane, the $y$ axis is oriented such that the incident electron beam (inclined by $70^{\circ}$ with respect to the surface normal) is in the $yz$ plane. The sizes of the maps, $10\times 10$ µm2 for sample 1 and $2\times 2$ µm2 for sample 2, are chosen to cover a sufficiently large number of dislocations, of the order of $10^{3}$, for a statistical analysis.
In a $2\times 2$ µm2 map of sample 0, just a single dislocation can be expected with a probability of 25%, compared to 500–700 dislocations in the maps of samples 1 and 2. So, sample 0 can be considered as dislocation free on the scale of the HR-EBSD measurements. Nevertheless, the range of the strain variation in Fig. 3(a) is comparable with that in samples 1 and 2. In contrast to these latter samples, the strain map does not reveal correlations between strains in neighboring measurement points, which is quantified below by calculation of the strain–strain correlations. Since the strain range for sample 0 in Fig. 3(a) is an order of magnitude larger than the expected sensitivity of the HR-EBSD measurements, the same measurement has been performed on a Si(001) wafer as presented in Fig. 3(d). The map also displays uncorrelated random noise, but the range of strain variations is of the order of $10^{-4}$, which is the expected sensitivity of HR-EBSD [25, 26] and an order of magnitude smaller than in all GaN samples under investigation. Whole sets of maps of all components of the strain and rotation tensors of Si(001) are presented in Fig. SM5 in the SM and show the variations in the same range of $10^{-4}$.
The map of $\varepsilon_{22}$ from sample 2 in Fig. 3(c) shows a notable difference between the $x$ and $y$ directions, with all features being extended in the $y$ direction. Figure SM4 shows a similar anisotropy in the maps of all strain and rotation components of this sample. We have performed the same measurement with the sample rotated by 90 around its normal and found that the direction of the extension in the maps is also rotated by 90 to the new $y$ direction. A possible anisotropy of the sample is thus ruled out and it is established that the features in the maps are extended along the direction of the tilt of the incident electron beam. We show below that this extension is explained by a strongly anisotropic resolution of the HR-EBSD measurements, and quantify this anisotropy.
[…]
Figure 4(a) presents by red and blue lines the probability distributions of the strain $\varepsilon_{22}$ obtained from the maps of samples 1 and 2 in Figs. 3(b) and 3(c). The dashed orange line is a similar distribution obtained from the map of the dislocation free GaN sample 0 in Fig. 3(a), while the thin black line is the respective distribution from the Si(001) sample in Fig. 3(d). A comparison of the widths of the strain distributions from dislocation free GaN and Si samples points to a material specific broadening for GaN. The possible origin of this broadening is discussed in Sec. 4 below. The contribution of dislocations in samples 1 and 2 to the widths of the strain distributions is only moderate.
The shear strains $\varepsilon_{13}$ and $\varepsilon_{23}$ are of separate interest. Since the information depth of EBSD is small (less than 20 nm), the conditions of the stress-free surface $\sigma_{i3}=0\,(i=1,2,3)$, are satisfied. One of these conditions, $\sigma_{33}=0$, is used in the processing of the Kikuchi patterns, since only the differences of the strains $\varepsilon_{11}-\varepsilon_{33}$ and $\varepsilon_{22}-\varepsilon_{33}$ can be found from the positions of the Kikuchi bands and this condition allows the three normal strain components to be determined separately [28]. The other two conditions $\sigma_{i3}=0\,(i=1,2)$ are not used in the processing of the Kikuchi patterns. Since $\sigma_{i3}=2c_{44}\varepsilon_{i3}\,(i=1,2)$ in a hexagonal crystal, the strains $\varepsilon_{13}$ and $\varepsilon_{23}$ must be zero at the free surface. Figure 4(b) shows the probability distributions of $\varepsilon_{13}$ obtained from the respective HR-EBSD maps of GaN samples 0, 1 and 2 as well as the silicon wafer. The probability distributions of the other shear strain component $\varepsilon_{23}$ are about two times narrower than those of $\varepsilon_{13}$. The maps of these components (shown in Figs. SM2–SM5 of the SM) exhibit uncorrelated random noise, as do all maps of the GaN sample 0 and the Si wafer. Figure 4(b) shows that all three GaN samples possess the same width of the $\varepsilon_{13}$ distribution, which is 16 times larger than the respective distribution for Si, and a factor of 2 narrower than the distribution of $\varepsilon_{22}$ from sample 0 in Fig. 4(a). Hence, the non-zero shear strains $\varepsilon_{13}$ and $\varepsilon_{23}$ are material specific. We postpone further discussion to Sec. 4.
Figure 5: Strain–strain correlation functions of (a,b) samples 1 and 2 and (c,d) their Monte Carlo simulations. The vertical red arrows point out to the screening distances taken as input of the Monte Carlo simulations (2 µm and 0.3 µm for samples 1 and 2 respectively).
Figures 5(a,b) and 6(a,b) present the strain-strain correlations calculated from the maps of $\varepsilon_{22}$ in Figs. 3(a,b) and from similar maps of other strain components presented in Figs. SM3 and SM4. We calculate the correlation functions as averages $\left\langle\varepsilon(\mathbf{r}_{1})\varepsilon(\mathbf{r}_{2})\right\rangle$ over the respective maps. Figure 5 shows correlations in $x$ and $y$ directions (the difference $\mathbf{r}_{1}-\mathbf{r}_{2}$ is directed either along $x$ or $y$), while Fig. 6 shows correlations as a function of the distance $r=\left|\mathbf{r}_{1}-\mathbf{r}_{2}\right|$ averaged over all orientations of $\mathbf{r}_{1}-\mathbf{r}_{2}$. Since the strain of a single dislocation decays with the distance as $\varepsilon(\mathbf{r})\propto r^{-1}$, we expect the correlations to decay as $\ln\left|\mathbf{r}_{1}-\mathbf{r}_{2}\right|$ as long as $\left|\mathbf{r}_{1}-\mathbf{r}_{2}\right|$ is less than the distance $R$, at which the dislocation strain is screened by surrounding dislocations, and to become zero at larger distances. We therefore plot the correlations in Figs. 5 and 6 as a function of the logarithm of the distance. The expected linear decay of the correlations in this scale and a kink from the logarithmic decay to zero correlations are clearly seen in Figs. 6(a,b) and allow the determination of the screening distances: these distances are 2 µm for sample 1 and 0.3 µm for sample 2. This finding is the main result of the present paper.
Figures 6(a,b) also show a drastic difference in the correlations between the in-plane strain components $\varepsilon_{11}$, $\varepsilon_{12}$, $\varepsilon_{22}$ and the shear strains $\varepsilon_{13}$, $\varepsilon_{23}$. The latter ones are expected to be zero due to zero stress conditions at the free surface but show distributions in Fig. 4(b) comparable in width to the in-plane strains. Their correlation functions in Figs. 6(a,b) are negligibly small compared to the correlations of the in-plane strains and quantify the visual impression of the respective maps in Figs. SM3 and SM4 as uncorrelated random noise.
One can see in Figs. 5 and 6 deviations from straight lines (i.e. from the logarithmic law) at separations smaller than about 200 nm: the curves are gradually flattened. One can also see in Fig. 5(b) a strong difference between correlations of $\varepsilon_{11}$ and $\varepsilon_{22}$ and a strong difference between the correlations of $\varepsilon_{22}$ in $x$ and $y$ directions. This plot thus quantifies the anisotropy of the map of $\varepsilon_{22}$ from sample 2 noted above. The respective plots for sample 1 in Fig. 5(a) do not show such a difference between correlations of $\varepsilon_{11}$ and $\varepsilon_{22}$ and between the $x$ and $y$ directions because of the larger step size employed for these maps.
3.3 Electron backscatter diffraction: Monte Carlo simulations
We simulate the maps of all strain and rotation components by the Monte Carlo method, as described in Sec. 2.4. We take the dislocation densities $5\times 10^{8}$ cm-2 and $1.8\times 10^{10}$ cm-2 for samples 1 and 2, as determined in Sec. 2.1, and the mean distances between dislocations in the pairs of 2 µm and 0.3 µm, respectively, as derived above from the plots in Fig. 6(a,b). We simulate maps having the same areas and number of pixels as in the experiment, with the aim to obtain the same level of statistical errors due to averaging over the same number of dislocations. The strains are integrated over a 20 nm thick layer at the surface, corresponding to the information depth of EBSD measurements. The simulated maps of $\varepsilon_{22}$ in Fig. 3(e,f), as well as the simulated maps of all strain and rotation components shown in Figs. SM6 and SM7 show a good qualitative agreement with the measured maps.
To make a quantitative comparison, we process the simulated maps to obtain the correlation functions in the same way as done for the measured maps. An agreement between simulations and experiment for sample 2 is reached when we introduce a strongly anisotropic resolution to the simulated maps. Specifically, the calculated strains are averaged over an area $20\times 200$ nm2, extended in the direction of inclination of the incident electron beam. Such a resolution allows to reach an agreement between the correlations in $x$ and $y$ directions, as seen in Figs. 5(b) and 5(d). The same resolution is imposed for sample 1 but does not give rise to a similar anisotropy of the maps, because the measurements are performed with a larger step size, to cover a larger surface area. As a result, both the experimental and the simulated correlation curves of sample 1 in Figs. 5(a) and 5(c) do not show the anisotropy observed for sample 2.
The measured strain distribution of sample 1 in Fig. 4(a) is slightly broader than the respective Monte Carlo simulation in Fig. 4(c). The additional broadening can be attributed to the apparent strain observed in sample 0 and discussed further in Sec. 4.4, which is independent of the dislocation strain.
Figure 4(d) presents the probability distributions of the strain component $\varepsilon_{13}$ obtained from the Monte Carlo simulated maps of samples 1 and 2 in Figs. SM6 and SM7, respectively. This strain is equal to zero at the surface because of the strain-free boundary conditions. For edge dislocations, $\varepsilon_{13}$ and $\varepsilon_{23}$ are also equal to zero in the bulk of the crystal far from the surface. The non-zero strains in Figs. 4(d) and SM6, SM7 result from integrating the three-dimensional strain relaxation field near the surface over the 20 nm-thick surface layer corresponding to the EBSD information depth. The widths of the probability distributions of $\varepsilon_{13}$ in the simulated maps in Fig. 4(d) are at least 5 times smaller than the respective experimental widths in Fig. 4(b). Hence, the strains $\varepsilon_{13}$, $\varepsilon_{23}$ in the experimental maps are not the strains near the surface. They look, similarly to all maps of the dislocation-free sample 0, like uncorrelated random noise and may have the same origin, which is discussed in Sec. 4.4.
The orientation-averaged correlation curves in Fig. 6(c,d), obtained from the Monte Carlo simulated maps, reveal the logarithmic decay of the strain correlations (straight lines in the coordinates of the figure) and a kink as a transition to zero correlations at distances exceeding the screening distance. The positions of the kinks coincide with the screening distances of 2 µm and 0.3 µm for samples 1 and 2 respectively, taken as an input of the Monte Carlo simulations. A gradual increase of correlations is seen at distances smaller than the resolution of 200 nm.
Figure 6: Orientation-averaged strain–strain correlation functions of (a,b) samples 1 and 2 and (c,d) their Monte Carlo simulations. The vertical red arrows point out to the screening distances taken as input of the Monte Carlo simulations (2 µm and 0.3 µm for samples 1 and 2 respectively).
The probability distributions and the correlation functions for the different rotation components, both the measured and the Monte-Carlo simulated ones, are shown in Figs. SM8 and SM9 of the SM. Edge threading dislocations dominate in sample 2 and give rise to only in-plane rotations $\omega_{12}$ in the bulk of the sample. However, strain relaxation at the free surface results in out-of-plane rotations $\omega_{13}$ and $\omega_{23}$. These rotations are smaller than $\omega_{12}$ but larger than the shear strains $\varepsilon_{13}$ and $\varepsilon_{23}$ in Fig. 4(d), since the out-of-plane rotations $\omega_{13}$ and $\omega_{23}$ for edge dislocations are not zero at the surface. The correlation functions of rotations presented in Fig. SM9 show the same logarithmic decay of correlations at distances smaller than $R$ and the absence of correlations at larger distances. The characteristic kink is seen at the same distances $R$ as in the strain correlations of the respective samples.
The mixed dislocations used to model sample 1 give rise to equal widths of probability distributions of the in-plane and out-of-plane rotations in the Monte Carlo simulations in Fig. SM8(c). The experimental distribution of the in-plane rotation $\omega_{12}$ in Fig. SM8(a) is broader than the distributions of $\omega_{13}$ and $\omega_{23}$. In order to achieve agreement, edge dislocations with a density of $1\times 10^{9}$ cm-2 would need to be added. However, that disagrees with the dislocation density obtained from the cathodoluminescence image in Fig. 1 and also with the close widths of the XRD profiles in symmetric and asymmetric Bragg reflections in Fig. 2(a), which point out to mixed dislocations as a main part of the dislocation ensemble in sample 1. Therefore, only mixed dislocations are included in the Monte Carlo simulations of sample 1.
As we noted in Sec. 2.3, the measured EBSD maps suffer from sample drift, which is estimated to be below 18%. While the drift has no effect on the strain probability distributions, it can potentially affect the correlations of the strain tensor components, which are the main focus of our study. However, 18% drift refers to the relative displacement of the opposite corners of a measured map. Since we are looking for correlations between points separated by distances smaller than the screening distance $R$ (the correlations are absent at larger distances anyway), and since the map sizes are chosen to be an order magnitude larger than $R$, the effect of drift on separations smaller than $R$ is negligible. This is quantified in Fig. SM10 in the SM. In this example, we artificially expand a Monte Carlo-simulated map by 15% and compare the autocorrelation functions of the strain maps before and after the stretching. The difference between the curves can be neglected.
To estimate a possible statistical error in the screening distance determination, we collect in a single plot the autocorrelation functions obtained in measurements at different areas of the samples. They are compared with repeated Monte Carlo simulations performed using the same parameters, but with different values for the seed of the random number generator. These plots are compared in Fig. SM11 of the SM. Similar scattering of the curves is observed in both the experiment and the simulations. Therefore, the limited number of dislocations in a map area (only 500–700) rather than possible sample inhomogeneity is a main factor limiting the accuracy of the screening length determination. This could potentially be improved by measuring maps from sample areas containing larger number of dislocations. From Fig. SM11, the accuracy of our determination of the screening distance is estimated to be about 20%.
4 Discussion
4.1 Strain distributions in XRD and HR-EBSD
Both XRD profiles and probability distributions obtained from HR-EBSD maps represent distributions of strains and rotations, as it was already pointed out and compared by Wilkinson et al. [27]. Measurements by these two methods are complimentary. The XRD intensity provides the probability distribution over several orders of magnitude down from the maximum, allowing the dislocation density $\varrho$ and the dislocation strain screening distance $R$ to be determined from a single profile. However, the XRD intensity is an average over the sample volume and potentially includes a contribution from strained regions far from the surface, in particular the film–substrate interface. HR-EBSD, in turn, provides the maps of strains and rotations at the surface. The statistics of the strain maps are limited by the number of pixels in the map. The maps can be used to determine not only the strain probability distributions, but also the spatial correlations of strains and rotations, which allows the screening distance $R$ to be determined directly as the distance at which the correlations become zero. A comparison of the dislocation correlations obtained by the two methods is the main aim of the present paper.
We have compared in Fig. 2 the measured XRD curves, their Monte Carlo simulations, and fits of both sets of curves by the formula proposed in Ref. [42]. A comparison of a curve simulated by the Monte Carlo method with the fit of this curve provides a check on the internal consistency of simulations and fits: dislocation densities obtained from fits of simulated curves (open symbols in Fig. 2(c)) coincide with the input values of the simulations with fairly little scatter.
The dislocation densities that are obtained in the Monte Carlo simulations and confirmed by the fits of the XRD curves are larger than the spot densities obtained from Fig. 1 by factors of 2 and 3 for samples 1 and 2, respectively. It is noteworthy that our previous studies [42, 54, 55] also yielded dislocation densities obtained from the XRD profiles that were several times larger than those determined by transmission electron microscopy. Since XRD collects intensity scattered from the whole volume of the GaN film, the additional broadening of XRD curves can be caused by strain originating from the film–substrate interface. However, calculations [55] show that misfit dislocations at the film–substrate interface modify the reciprocal space maps but, for thick films, have little effect on the double-crystal curves that are measured in the present work.
4.2 Screening distances
The screening of dislocation strains by surrounding dislocations provides a transition from the Gaussian central part of the XRD profile to the $I(q)\propto q^{-3}$ tails. The screening distance $R$ is a second parameter, in addition to the dislocation density $\varrho$, that is included in the fit of the XRD profiles [42]. The screening distances obtained by fitting the experimental curves (full symbols in Fig. 2(d)) agree with those obtained from correlations in HR-EBSD maps (horizontal dashed lines in Fig. 2(d)) within a factor of 2, with a rather large scatter between the values obtained from different reflections.
Plots of the autocorrelation functions of the components of the strain and rotation tensors measured by HR-EBSD provide a more robust tool for determining the screening distance $R$. This distance is clearly seen as a kink in the plot of the correlation function in dependence of the logarithm of the separation, and is found to be the same for all strain and rotation components. We find screening distances $R$ of 2 µm and 0.3 µm for the GaN films with threading dislocation densities $\varrho$ of $5\times 10^{8}$ cm-2 and $1.8\times 10^{10}$ cm-2, respectively. In both cases, the number of dislocations that provide screening of the strain of a given dislocation $M=R\varrho^{1/2}$ is about 4. This means that the dislocation strain field is very effectively screened by only a few surrounding dislocations. To obtain these screening distances, we measure the strain and rotation maps over areas with linear sizes that are large compared to $R$ using steps that are small compared to $R$. Different map sizes and steps are chosen depending on the dislocation density. We note that very close numbers, 4 to 6 dislocations providing screening of the dislocation strain fields, are obtained for plastically deformed metals [56] by an XRD line profile analysis method that is close to our approach for GaN films.
The screening distances are obtained directly by calculating the correlation functions from the measured maps and do not require simulations. The same analysis can be carried out for more complicated dislocation distributions, when a simulation of the strain maps is not as straightforward as in our case. We have performed simulations of the maps to better understand them, but this is not necessary in every case where the screening distance is of interest.
4.3 Geometrically necessary dislocations
Maps of the strain and rotation components are commonly used in HR-EBSD studies to determine the densities of geometrically necessary dislocations (GND) [57, 58, 59, 60, 61, 62]. The GND density corresponds to the minimum density of dislocations that is required to provide a net effect that is of interest [63, 64]. It is therefore dependent on the scale of that effect and implies an appropriate spatial averaging of the strains and rotations at smaller scales. The papers cited above use the concept of GND to describe boundaries between misoriented domains or plastic bending of a crystal. The GND value is then a minimum density of dislocations that provides a small angle boundary or the curvature of the crystal lattice. In these cases, strain variations can be neglected compared to the lattice rotations, and the GND can be obtained from the differences in rotations between neighboring points in the measured maps.
Figure 7: (a) Side view of the electron trajectories in GaN simulated by CASINO, (b) top view of the exit points of the trajectories of backscattered electrons and (c,d) distributions of the exit points in projections on the $x$ and $y$ axes. Dashed lines are the distributions of the exit points of the electrons that lost less than 10% of the initial energy. The incident electron beam (shown by a red arrow) of energy 15 keV is inclined by 70 along the $y$ axis.
Threading dislocations in GaN films are an example of a system where the concept of GND requires a different consideration. The screening distances $R$ that we have found provide a scale for the determination of the GND. Counting dislocations on a smaller scale gives the density of statistically stored dislocations, while the strain averaged over distances exceeding $R$ can be treated as due to the GND. The GND density for threading dislocations in GaN films is zero, since the dislocations screen each other’s strain fields and do not provide a net long-range effect. More specifically, the strains and rotations are zero after averaging over the areas exceeding the screening distance $R$ in the respective sample. To obtain this zero average, the step size of a measurement has to be small compared to $R$. In the studies cited above, the GND density is obtained from the differences in rotations between neighboring points without an average over the points that would be measured in between. In our samples, strain and rotation maps measured with a step larger than $R$ would only show an uncorrelated random distribution with the same probability density as our strain maps. The difference between neighboring measured points would not contain any feasible information if the step of the measurements is chosen larger than $R$. On the other end, a measurement with a step small compared to $R$ would provide a local dislocation density that is needed to generate the particular strain and rotation maps. This latter treatment of GND is applied in Ref. [28] to a very similar case of threading dislocations in (In,Al)N films.
4.4 Spatial resolution of HR-EBSD maps
We find that the correlation functions deviate from the $\propto\ln r$ law at small distances and become more gradual. Also, the HR-EBSD maps of sample 2 with high dislocation density measured with a step between the measuring points of only 20 nm, show a clear distinction between the $x$ and $y$ directions. A comparison of the correlations in the $x$ and $y$ directions and Monte Carlo simulations allow us to estimate the spatial resolution of the HR-EBSD maps to be $20\times 200$ nm2, with the worse resolution along the direction of the tilt of the incident electron beam. In metals, grain boundaries are used for a direct measurement of the resolution, which depends on the material and the accelerating voltage (we refer to the papers [65, 66] which also provide literature reviews). The resolutions in $x$ and $y$ directions are obtained by appropriately orienting the grain boundary with respect to the direction of the electron beam. The resolution that we obtain is comparable to that obtained in relatively light metals but shows a larger anisotropy between the $x$ and $y$ directions.
For additional insight, we simulate the trajectories of the electrons in GaN using the free software CASINO [67]. Figure 7(a) is a side view on the electron trajectories. Red lines are trajectories of the backscattered electrons. A top view of their exit points is shown in Fig. 7(b). Their spatial distribution is very asymmetric with respect to the $x$ and $y$ directions. The projections of the outcrop distributions on the $x$ and $y$ axes in Figs. 7(c,d) are even broader than the resolution that we obtain from the simulations of the strain and rotation maps. Since low-loss electrons provide the major contribution to HR-EBSD patterns [68], we plot by dotted lines in Figs. 7(c,d) the distributions of electrons that lost less than 10% of their initial energy of 15 keV. The widths of these distributions are consistent with the resolution that we found.
The HR-EBSD maps of the dislocation-free GaN sample 0 show only uncorrelated random noise for all strain and rotation components. The widths of these distributions are an order of magnitude larger compared to Si(001). We also note that the XRD curves of this sample, albeit narrow compared to the other samples, are significantly broader than expected for its dislocation density [31]. The HR-EBSD measurements on GaN sample 0 were performed before and after surface cleaning, which had a negligible effect on the strain distributions. Hence, neither surface roughness nor contamination do appear to be responsible for the apparent strain in this sample.
In the absence of dislocations, the remaining source of strain are point defects, including impurities as well as native defects. The density of native defects in GaN is orders of magnitude larger than in Si. The density of Ga vacancies alone in GaN ranges from $10^{16}$$10^{18}$ cm-3 [69, 70, 71], as compared to a vacancy density below $1\times 10^{14}$ cm-3 in Si [72, 73]. To examine the effect of point defects, we have performed simulations of their strain maps for a spatially random distribution of dilatation centers with the excess volume equal to the atomic volume. Even for a density of $10^{18}$ cm-3, the strain produced by these centers is too low to account for the experimentally measured strain in Fig. 3(a). Moreover, the simulated strain maps exhibit the same anisotropy as seen in Fig. 3(c) caused by the inclination of the incident electron beam.
The distributions of the shear strains $\varepsilon_{13}$ and $\varepsilon_{23}$ coincide in all three GaN samples and are therefore independent of the dislocation density. These strains have to be zero due to the stress free boundary conditions. Of the three zero stress conditions at the surface $\sigma_{i3}=0$ ($i=1,2,3)$, one condition, $\sigma_{33}=0$, is used in the processing of the Kikuchi patterns. The other two conditions reduce to $\varepsilon_{13}=0$ and $\varepsilon_{23}=0$ for the (0001) surface of a hexagonal crystal. The distributions of these components thus provide an estimate of the sensitivity limit to dislocation strains. We note that the maps of these components do not show any correlations in the crystals with dislocations, in contrast to other strain and rotation components that develop correlations. Thus, in the case where a dislocation free crystal is not available as a reference, the distributions of these shear strain components can provide an estimate of the accuracy of the other strain and rotation maps.
5 Summary
The screening of the dislocation strains by the surrounding dislocations is revealed by both fits of the XRD curves in reciprocal space and the real space correlations in the HR-EBSD strain maps. Our study is performed using a well-defined array of threading dislocations in GaN(0001) as an example. It allowed Monte Carlo simulations of XRD curves and HR-EBSD maps within one and the same model of dislocation distributions and its quantitative comparison with the experiment. The simulations provide a basis for determining of the screening distances from the HR-EBSD maps as the positions of the kinks in the strain autocorrelations plotted against the logarithm of the distance. Determining the screening distance in this way does not require any simulations or fits and provides a robust tool also for more complicated dislocation ensembles, where simulations would be more difficult.
In the XRD, dislocation densities and screening distances are obtained by fitting the diffraction profiles. We find that XRD overestimates the dislocation densities, presumably due to contributions from the strained region near the film–substrate interface. The screening distances obtained by XRD from different reflections possess larger scattering and are consistent with those obtained by HR-EBSD within a factor of two.
Screening distances of 2 µm and 0.3 µm are obtained for the GaN films with threading dislocation densities of $5\times 10^{8}$ cm-2 and $1.8\times 10^{10}$ cm-2, respectively. This indicates that the dislocation strain is screened by only four neighboring dislocations in both samples. The strain maps from the high dislocation density sample measured with a small step size show a large anisotropy in the resolution of the HR-EBSD measurements. The resolution is estimated to be $20\times 200$ nm2, with the worse resolution in the direction corresponding to the inclination of the incident electron beam.
Acknowledgments
The authors thank Bernd Jenichen for XRD measurements on sample 2 made for our former study [31], Carsten Richter for useful discussions, Jingxuan Kang for cleaning sample 0 and Moritz Hansemann for AFM measurements of that sample, as well as Achim Trampert for a critical reading of the manuscript. Arno Winkelmann and Graham Meaden read the preprint of the present paper and analyzed our HR-EBSD data with their software. We are grateful to them for their efforts and comments.
Supplementary material to the paper
Dislocation correlations in GaN epitaxial films
revealed by EBSD and XRD
by
Vladimir M. Kaganer, Domenik Spallek, Philipp John, Oliver Brandt, and Jonas Lähnemann
SM  1 Kikuchi patterns
Figure SM1: Kikuchi patterns of the GaN samples 0, 1, 2 and the Si wafer.
SM  2 Strain and rotation maps
Figure SM2: Components of the strain and rotation tensors of sample 0 (dislocation free GaN sample).
Figure SM3: Components of the strain and rotation tensors of sample 1.
Figure SM4: Components of the strain and rotation tensors of sample 2.
Figure SM5: Components of the strain and rotation tensors of a silicon wafer. Note different ranges of the values in comparison to Fig. SM2.
Figure SM6: Monte Carlo simulation of the strain and rotation tensors of sample 1.
Figure SM7: Monte Carlo simulation of the strain and rotation tensors of sample 2.
SM  3 Probability distributions and correlation functions of rotations
Figure SM8: Probability distributions of rotations (a,b) in the measured maps and (c,d) in Monte Carlo simulations.
Figure SM9: Orientation-averaged autocorrelation functions of the rotations (a,b) in the measured maps and (c,d) in Monte Carlo simulations.
SM  4 Effect of drift
Figure SM10 simulates the effect of drift on the correlation functions of strain. Figure SM10(a) is the Monte Carlo simulated map of the strain component $\varepsilon_{22}$ of sample 2. The same map is presented in Fig. 3(f) in the main text and in Fig. SM7. In Fig. SM10(b), a 15% drift in $y$-direction is simulated, assuming that the effect of drift is accumulated linearly: each pixel of the map is shifted vertically proportional to the time between the measurement of that pixel and the first pixel, so that the last pixel is shifted by 15%. Note that the drift effect is negligible at the top of the map and continuously increases to the bottom. The bottom part of the expanded map is cut out, to have the same area as in the initial map. Figure SM10(c) compares the autocorrelation functions obtained from these two maps. Since the correlations tend to zero at distances exceeding 0.2 µm, which is 0.1 of the map size, only correlations of strains in closely spaced points are of interest. As a result, the effect of drift on the autocorrelation function in Fig. SM10(c) is negligibly small.
Figure SM10: (a) Monte Carlo simulated map of the strain $\varepsilon_{22}$ of sample 2 (identical to Fig. 3(f) in the main text and Fig. SM7), (b) the same map with a 15% drift in $y$-direction added, and (c) autocorrelation functions obtained from these maps.
SM  5 Repeated measurements and Monte Carlo simulations of EBSD maps
Figure SM11: Autocorrelation functions of the strain component $\varepsilon_{22}$ obtained (a,b) in the HR-EBSD study of samples 1 and 2 and (c,d) by repeating Monte Carlo simulations of HR-EBSD maps with all parameters kept the same as in Figs. 3–6. Blue lines are the calculations in Fig. 6. Black lines in the experimental plots are due to the measurements in different places of the respective samples, while black lines in the Monte Carlo modeling are obtained by changing the seed of the random number generator. The vertical red arrows point out to the screening distances taken on input of the Monte Carlo simulations (2 µm and 0.3 µm for samples 1 and 2 respectively).